OpenStudy (anonymous):

2x + 90 + (x+15) = 180 ( what is the value of x ?)

jimthompson5910 (jim_thompson5910):

2x+90+(x+15)=180 2x+90+x+15=180 3x+105=180 3x=180-105 3x=75 x=75/3 x=25

OpenStudy (anonymous):

2x+x+90+15=180 3x+105=180 3x=75 x=25